OpenStudy (anonymous):

The sum of one half of Abdul's age and 44 is equal to his father's age. If Abdul's father is 54 years old, how old is Abdul?

OpenStudy (anonymous):

you transform that into an equation abduls age=n so \[\frac{ n }{ 2 }+44=54, \frac{ n }{ 2 }=54-44=10,n=10*2=20\]

OpenStudy (anonymous):

thank you i am your fan :D keep up the great work

OpenStudy (anonymous):

:D thank you very much, I'm glad to be of help